# Nonparametric Estimation of Matching Efficiency and Mismatch in Labor Markets via Public Employment Security Offices in Japan, 1972-2024

### v1

@@ Abstract
I examine changes in matching efficiency and elasticities in Japan’s labor market via Hello Work for unemployed workers from January 1972 to April 2024 using a nonparametric identification approach by Lange and Papageorgiou (2020). I find a declining trend in matching efficiency, consistent with decreasing job and worker finding rates. The implied match elasticity with respect to unemployment is 0.5-0.9, whereas the implied match elasticity with respect to vacancies is 0.1-0.4. Decomposing aggregate data into full-time and part-time ones, I find that the sharp decline of matching efficiency after 2015 shown in the aggregate trend is driven by the decline of both full-time and part-time ones. Second, I extend the mismatch index proposed by Şahin et al. (2014) to the nonparametric version and develop the computational methodology. I find that the mismatch across job categories is more severe than across prefectures and the original Cobb-Douglas mismatch index is underestimated.
@@ 1 Introduction
The main empirical findings of this paper are twofold. First, matching efficiency (normalized to 1972) in the labor market via Hello Work shows a declining trend with notable fluctuations, which is consistent with downward trends of job and worker finding rates. This might be because matching opportunity out of the government-operated platform has increased. Implied match elasticity with respect to unemployment is 0.5-0.9, which is comparable to previous worldwide findings such as Petrongolo and Pissarides (2001) (the range in 0.5-0.7) and Japanese studies. On the other hand, implied match elasticity with respect to vacancies is 0.1-0.4, which is comparable to Lange and Papageorgiou (2020) (0.15-0.3) and Japanese studies.
Second, substantial mismatch are observed in both prefecture-level and job-category-level and the trend is upward with fluctuations. The mismatch across prefectures is 0.3 on average, whereas the mismatch across job categories is 0.4-0.65, larger than previous studies such as Shibata (2020) and Higashi and Sasaki (2023). The nonparametric mismatch index highlights the downward bias of the original Cobb-Douglas mismatch index through Monte Carlo simulation and empirical application.
@@ 1.1 Related Literature
Second, this paper contributes to the literature on mismatch Şahin et al. (2014). The approach is widely used in a variety of contexts, for example, United Kingdom in Patterson et al. (2016). In the context of Japanese labor markets, several papers estimate mismatch at geographical- and job-category-level using Cobb-Douglas specification (Shibata 2013, 2020, Kawata et al. 2016, Higashi 2018, Kawata 2019, Higashi 2020, 2021, Higashi and Sasaki 2023). This paper points out that Cobb-Douglas specification cancels out time variations of matching efficiency which induce bias of the mismatch index. Then, this paper proposes the nonparametric version of the mismatch index, and applies a novel computational method known as Mathematical Programming with Equilibrium Constraint (MPEC) proposed by Su and Judd (2012), Dubé et al. (2012) to calculate the geographical- and job-category-level mismatch index.
@@ 3.1 Matching efficiency and elasticities
Finally, for calculating matching elasticities, I run a LASSO regression projecting hires on the original and squared numbers of vacancies and unemployed interacted with implied matching efficiency. The estimates approximate the derivatives of the matching function with respect to vacancies and unemployed interacted with implied matching efficiency, that is, an estimate of the elasticity of the matching function. [^6] The matching elasticity with respect to unemployed $\frac{d\log m(AU,V)}{d\log U}=\frac{dm(AU,V)}{dU}\frac{U}{H}=\frac{dm(AU,V)}{dAU}\frac{dAU}{dU}\frac{U}{H}=\frac{dm(AU,V)}{dAU}\frac{AU}{H}=\frac{d\log m(AU,V)}{d\log AU}$ is obtained from the regression coefficient of $H$ on $AU$ and multiplying it by $\frac{AU}{H}$. Concretely, we approximate $m$ by the second order polynomial $m=\beta_{1}(AU)+\beta_{2}(AU)V+\beta_{3}V+\beta_{4}(AU)^{2}+\beta_{5}V^{2}$ and get $\frac{dm(AU,V)}{dAU}=\beta_{1}+\beta_{2}V+2\beta_{4}(AU)$ and $\frac{d\log m}{d\log U}=\frac{d\log m}{d\log AU}=(\beta_{1}+\beta_{2}V+2\beta_{4}(AU))\frac{U}{H}$ and $\frac{d\log m}{d\log V}=(\beta_{2}(AU)+\beta_{3}+2\beta_{5}V)\frac{V}{H}$. For calculating mismatch, we need $\frac{dm}{dU}=\frac{d\log m(AU,V)}{d\log AU}\frac{H}{U}=\beta_{1}+\beta_{2}V+2\beta_{4}(AU)$.
@@ 4 Data
The three types of number of job openings, job seekers, and successful job placements are reported in the data. The first is the number for full-time and part-time workers and jobs. The second is the number for full-time workers and jobs. The third is the number for part-time workers and jobs. The first is the sum of the second and third. Using the three types of data, I can decompose the labor market features into full-time and part-time ones.
Second, for estimating nonparametrically worker-job mismatch across regions and occupation categories, I use these submarket data in 2012-2024. The period limitation shared with Kawata (2019) and Higashi (2018) is due to the revision of job classifications before and after 2012, which makes it difficult to accurately connect the data. Finally, I use all 47 prefectures and 67 job categories in each month as submarkets.
@@ 5.1 Aggregate trends in 1972-2023
Figures 2 (a)-(d) provide monthly data patterns of unemployed individuals, vacancies, labor market tightness ($V/U$), hires, the ($U$,$V$) relationship, and job and worker finding rates ($M/U$ and $M/V$) in aggregate labor markets. In summary, the numbers of unemployed individuals and vacancies increase with fluctuations, corresponding with market tightness, while the number of hires shows a noticeable decline around the late 1970s, peaks and troughs from the mid-1980s to the late 1990s, peaks around mid-2000s and 2010, and a sharp decline towards recent years. Although the ($U$,$V$) relationship does not corresponds with Beveridge curve because these numbers are not divided by the corresponding number of labor force, the prominent rightward shifts are consistent with the previous studies (Elsby et al. 2015).
@@
Figures 2 (e) and (f) present the estimation results of the matching function along with matching efficiency and elasticities ($\frac{d\ln M}{d\ln AU}$ and $\frac{d\ln M}{d\ln V}$). Notably, matching efficiency (normalized to 1972) shows a declining trend with notable fluctuations, which is consistent with the downward trends of job and worker finding rates. In particular, the significant decline after 2015 is remarkable. This seems due to an increase in matching opportunities outside of the government-operated platform.
@@
The implied match elasticity with respect to unemployment is 0.5-0.9, which is comparable to previous worldwide findings such as Petrongolo and Pissarides (2001) (range: 0.5-0.7) and Japanese studies such as Higashi (2018) (0.38 for 2000-2014 monthly), Kawata (2019) (0.48 for 2012-2017 prefecture-month-level), Kano and Ohta (2005) (0.56 for 1972-1999 prefecture-year-level), Sasaki (2007) (about 0.6 for 1998-2007 prefecture-quarter-level), and Kambayashi and Ueno (2006) (about 0.8 for 1996-2001 prefecture-month-level). [^8] For reference, Petrongolo and Pissarides (2001) summarize early aggregate studies in many countries based on a Cobb-Douglas matching function with the flow of hires on the left-hand side and the stock of unemployment and job vacancies on the right-hand side. In short, match elasticity with respect to unemployment is in the range of 0.5–0.7. Bernstein et al. (2022) review the recent empirical literature that estimates the matching elasticity in the U.S. by four types of specification; Cobb-Douglas, Cobb-Douglas with endogeneity correction, a constant elasticity of substitution (CES), and nonparametric (Lange and Papageorgiou 2020). The broad range of estimates is partly due to variations in data choices and periods. Higher estimates are typically derived from JOLTS data (Borowczyk-Martins et al. 2013, Şahin et al. 2014), while lower estimates are often obtained from CPS flows data (Barnichon and Figura 2015) or job-category-level hires from CPS data (Şahin et al. 2014). On the other hand, the implied match elasticity with respect to vacancies is 0.1-0.4, which is comparable to Lange and Papageorgiou (2020) (range: 0.15-0.3) and Japanese studies such as Higashi (2018) (0.24 for 2000-2014 monthly), Kawata (2019) (0.52 for 2012-2017 prefecture-month-level), Kano and Ohta (2005) (0.3 for 1972-1999 prefecture-year-level), Sasaki (2007) (about 0.2 for 1998-2007 prefecture-quarter-level), and Kambayashi and Ueno (2006) (about 0.3 for 1996-2001 prefecture-month-level).
@@
5.2 Decomposition of part-time and full time trends in 1972-2023
@@
Given estimation results in each dataset, Figure 4 (e) depicts the trend of implied matching efficiency. The downward trends are common. Full-time matching efficiency fluctuates from 80% to 300%, whereas part-time matching efficiency remains stable at 60% compared to January 1972. This concludes that the sharp decline of matching efficiency after COVID-19 shown in the aggregate trend is driven by the decline of full-time matching efficiency.
Figure 4 (f) depicts the trends of matching elasticities. For matching elasticity with respect to unemployed, the full-time trend exhibits significant fluctuations, especially before the 1980s, ranging from 0.2 to around 1.0, indicating varying sensitivity of matching efficiency to changes in unemployment over time. In contrast, the part-time trend peaks, ranging from 0.9 to 1.0 after 2020, which may indicate the post-COVID effect on the elasticity. For matching elasticity with respect to vacancies, the full-time trend shows a relatively stable trend, starting around 0.2, with slight increases in recent years, whereas the part-time trend maintains lower and more stable values around 0.1 to 0.3, suggesting that vacancies have a less pronounced effect on matching efficiency in the part-time labor market.
Figure 4 illustrates some correlation patterns between matching efficiency and market structure variables corresponding with Figures 2 (g)-(h). I find the stronger and more consistent relationships in the full-time labor market. This confirms that positive correlations between efficiency and market structure driven mainly by full-time markets.
@@ 5.3 Prefecture-level aggregate results in 2012-2024
Figure 5 displays the month-prefecture level matching efficiency, normalized to 2013 January Tokyo, across different regions of Japan from 2012 to 2024. Each panel shows the efficiency trends within a specific region, highlighting notable regional variations. For instance, the Tohoku region (panel b) and Chubu region (panel d), which are rural areas, exhibit relatively high volatility and generally higher efficiency values compared to other regions, suggesting greater variability in matching efficiency. In contrast, regions such as Kanto (panel c) and Kansai (panel e) show more stable and lower efficiency levels. The Tohoku area is about twice as efficient as the Tokyo area. These patterns suggest that regional economic factors and labor market conditions significantly influence matching efficiency across Japan.
@@
5.4 Job category level results in 2012-2024
Figure 8 illustrates significant disparities in matching efficiency across job categories, normalized to 2013 January for clerical jobs. Primary and secondary industries show relatively higher efficiency than tertiary industries. For example, categories like ”Agriculture, Forestry, and Fishing” (panel f) and ”Construction and Mining” (panel i) in primary and secondary industries exhibit high volatility and generally higher efficiency levels. In contrast, job categories such as ”Managerial” (panel a) and ”Service” (panel e) display more stable and lower efficiency values. These variations highlight the diverse nature of labor market dynamics across different sectors, influenced by factors such as skill specialization, economic cycles, and industry-specific demand fluctuations.
Figure 8: Month-job category level matching efficiency results 2012-2024
Figures 9 and 10 illustrate matching elasticities with respect to unemployed and vacancies in each job category in 2012-2024. As in the prefecture-level results, there is significant regional heterogeneity within the specific job area. The matching elasticities with respect to unemployed and vacancies fluctuates with a range generally between 0.1 and 6.0, which is larger than prefecture-level results. This indicates that matching heterogeneity depends on the definition of markets and matters in job-category level rather than geographical level.
Figure 9: Month-job category level matching elasticity with respect to unemployed 2012-2024
Figure 10: Month-job category level matching elasticity with respect to vacancies 2012-2024

### v2

@@ Abstract
I examine changes in matching efficiency and elasticities in Japan’s labor market via Hello Work for unemployed workers from January 1972 to April 2024 using a nonparametric identification approach by Lange and Papageorgiou (2020). I find a declining trend in matching efficiency, consistent with decreasing job and worker finding rates. The implied match elasticity with respect to unemployment is 0.5-0.9, whereas the implied match elasticity with respect to vacancies varies between -0.4 and 0.4. Decomposing aggregate data into full-time and part-time ones, I find that the sharp decline of matching efficiency after 2015 shown in the aggregate trend is driven by the decline of both full-time and part-time ones. Second, I extend the mismatch index proposed by Şahin et al. (2014) to the nonparametric version and develop the computational methodology. I find that the mismatch across occupations is more severe than across prefectures and the original Cobb-Douglas mismatch index is underestimated.
@@ 1 Introduction
The main empirical findings of this paper are twofold. First, matching efficiency (normalized to 1972) in the labor market via Hello Work shows a declining trend with notable fluctuations, which is consistent with downward trends of job and worker finding rates. This might be because matching opportunity out of the government-operated platform has increased. Implied match elasticity with respect to unemployment is 0.5-0.9, which is comparable to previous worldwide findings such as Petrongolo and Pissarides (2001) (the range in 0.5-0.7) and Japanese studies. On the other hand, implied match elasticity with respect to vacancies varies between -0.4 and 0.4, which is more volatile than Lange and Papageorgiou (2020) (0.15-0.3) and Japanese studies (0.15-0.6).
Second, substantial mismatch are observed in both prefecture-level and occupation-level and the trend is upward with fluctuations. The mismatch across prefectures is 0.3 on average, whereas the mismatch across occupations is 0.4-0.65, larger than previous studies such as Shibata (2020) and Higashi and Sasaki (2023). The nonparametric mismatch index highlights the downward bias of the original Cobb-Douglas mismatch index through Monte Carlo simulation and empirical application.
@@ 1.1 Related Literature
Second, this paper contributes to the literature on mismatch (Şahin et al. 2014). The approach is widely used in a variety of contexts, for example, United Kingdom in Patterson et al. (2016). In the context of Japanese labor markets, several papers estimate mismatch at geographical- and occupation-level using Cobb-Douglas specification (Shibata (2013, 2020), Kawata et al. (2016), Higashi (2018), Kawata (2019), Higashi (2020, 2021), Higashi and Sasaki (2023)). This paper points out that Cobb-Douglas specification cancels out time variations of matching efficiency which induce bias of the mismatch index. Then, this paper proposes the nonparametric version of the mismatch index, and applies a novel computational method known as Mathematical Programming with Equilibrium Constraint (MPEC) proposed by Su and Judd (2012) and Dubé et al. (2012) to calculate the geographical- and occupation-level mismatch index.
@@ 3.1 Matching efficiency and elasticities
Finally, for calculating matching elasticities, I run a LASSO regression projecting hires on the original and squared numbers of vacancies and unemployed interacted with implied matching efficiency. The estimates approximate the derivatives of the matching function with respect to vacancies and unemployed interacted with implied matching efficiency, that is, an estimate of the elasticity of the matching function. [^6] The matching elasticity with respect to unemployed $\frac{d\log m(AU,V)}{d\log U}=\frac{dm(AU,V)}{dU}\frac{U}{H}=\frac{dm(AU,V)}{dAU}\frac{dAU}{dU}\frac{U}{H}=\frac{dm(AU,V)}{dAU}\frac{AU}{H}=\frac{d\log m(AU,V)}{d\log AU}$ is obtained from the regression coefficient of $H$ on $AU$ and multiplying it by $\frac{AU}{H}$. Concretely, we approximate $m$ by the second order polynomial $m=\beta_{1}(AU)+\beta_{2}(AU)V+\beta_{3}V+\beta_{4}(AU)^{2}+\beta_{5}V^{2}$ and get $\frac{dm(AU,V)}{dAU}=\beta_{1}+\beta_{2}V+2\beta_{4}(AU)$ and $\frac{d\log m}{d\log U}=\frac{d\log m}{d\log AU}=(\beta_{1}+\beta_{2}V+2\beta_{4}(AU))\frac{U}{H}$ and $\frac{d\log m}{d\log V}=(\beta_{2}(AU)+\beta_{3}+2\beta_{5}V)\frac{V}{H}$. For calculating mismatch, we need $\frac{dm}{dU}=\frac{d\log m(AU,V)}{d\log AU}\frac{H}{U}$ at $U=U^{*}$ and $H=H^{*}=m(AU^{*},V)$.
@@ 3.2 Mismatch
4 Monte Carlo simulation
  • How large is the bias of matching efficiency from the Cobb Douglass specification?
  • How large is the bias of mismatch from the Cobb Douglass specification?
  • Mismatch literature:
    • mismatch Shimer (2007)
    • mismatch and large firms (Eeckhout 2018)
    • Classical Jackman and Roper (1987)
@@ 5 Data
We study the labor markets for full-time and part-time workers and jobs. In the Hello Work system, part-time workers are defined as those who work fewer hours than regular employees at the same establishment. MHLW classifies part-time workers into two categories: (1) regular part-time workers, who have contracts for an indefinite period or a period exceeding four months, and (2) temporary part-time workers, whose contracts last between one and four months or whose employment is fixed and typically tied to seasonal demand. While in many countries, part-time and full-time jobs are distinguished by the number of hours worked, in Japan, these terms often imply additional differences, such as variations in responsibilities, benefits, and flexibility. Part-time workers generally have fewer responsibilities and benefits but enjoy more flexible working hours compared to full-time employees. Full-time workers are those who are not part-time workers.
The three types of number of job openings, job seekers, and successful job placements are reported in the data. The first is the number for full-time and part-time workers and jobs. The second is the number for full-time workers and jobs. The third is the number for part-time workers and jobs. The first is the sum of the second and third. Note that data on unemployed, vacancies, and hires exclude workers newly graduated from college, and are used to calculate the unemployment rate for Japan. Using the three types of data, I can decompose the labor market features into full-time and part-time ones.
Second, to estimate worker-job mismatch nonparametrically across regions and occupation categories, I use submarket data from January 2012 to March 2023 for prefecture-level analysis, and up to March 2024 for occupation category-level analysis. The time frame, consistent with Kawata (2019) and Higashi (2018), is constrained by the revision of job classifications before and after 2012, which complicates accurate data connection. Finally, I analyze all 47 prefectures and 67 occupations, treating each as a submarket on a monthly basis.
@@ 6.1 Aggregate trends in 1972-2023
Figures 2 (a)-(d) provide monthly data patterns of unemployed individuals, vacancies, labor market tightness ($V/U$), hires, the ($U$,$V$) relationship, and job and worker finding rates ($M/U$ and $M/V$) in aggregate labor markets. [^8] Worker finding rates are also known as vacancy yields. Davis et al. (2013) define vacancy yields in month $t$ as the number of hires in month $t\left(h_{t}\right)$ divided by the number of vacancies in the previous month $\left(v_{t-1}\right)$ instead of the same month $\left(v_{t}\right)$, the two series are very similar. In summary, the numbers of unemployed individuals and vacancies increase with fluctuations, corresponding with market tightness, while the number of hires shows a noticeable decline around the late 1970s, peaks and troughs from the mid-1980s to the late 1990s, peaks around mid-2000s and 2010, and a sharp decline towards recent years. Although the ($U$,$V$) relationship does not corresponds with Beveridge curve because these numbers are not divided by the corresponding number of labor force, the prominent rightward shifts are consistent with the U.S. studies (Elsby et al. 2015).
@@
Figures 2 (e) and (f) present the estimation results of the matching function along with matching efficiency and elasticities ($\frac{d\ln M}{d\ln AU}$ and $\frac{d\ln M}{d\ln V}$). Notably, matching efficiency (normalized to 1972) shows a declining trend with notable fluctuations, which is consistent with the downward trends of job and worker finding rates. In particular, the significant decline after 2015 is remarkable. This seems due to an increase in matching opportunities outside of the government-operated platform, which is discussed in Otani (2024) using proprietary on-the-job search platform data.
@@
The implied match elasticity with respect to unemployment is 0.5-0.9, which is comparable to previous worldwide findings such as Petrongolo and Pissarides (2001) (range: 0.5-0.7) and Japanese studies such as Higashi (2018) (0.38 for 2000-2014 monthly), Kawata (2019) (0.48 for 2012-2017 prefecture-month-level), Kano and Ohta (2005) (0.56 for 1972-1999 prefecture-year-level), Sasaki (2007) (about 0.6 for 1998-2007 prefecture-quarter-level), and Kambayashi and Ueno (2006) (about 0.8 for 1996-2001 prefecture-month-level). [^9] For reference, Petrongolo and Pissarides (2001) summarize early aggregate studies in many countries based on a Cobb-Douglas matching function with the flow of hires on the left-hand side and the stock of unemployment and job vacancies on the right-hand side. In short, match elasticity with respect to unemployment is in the range of 0.5–0.7. Bernstein et al. (2022) review the recent empirical literature that estimates the matching elasticity in the U.S. by four types of specification; Cobb-Douglas, Cobb-Douglas with endogeneity correction, a constant elasticity of substitution (CES), and nonparametric (Lange and Papageorgiou 2020). The broad range of estimates is partly due to variations in data choices and periods. Higher estimates are typically derived from JOLTS data (Borowczyk-Martins et al. 2013, Şahin et al. 2014), while lower estimates are often obtained from CPS flows data (Barnichon and Figura 2015) or occupation-level hires from CPS data (Şahin et al. 2014). On the other hand, the implied match elasticity with respect to vacancies is 0.1-0.4 before 2015, which is comparable to Lange and Papageorgiou (2020) (range: 0.15-0.3) and Japanese studies such as Higashi (2018) (0.24 for 2000-2014 monthly), Kawata (2019) (0.52 for 2012-2017 prefecture-month-level), Kano and Ohta (2005) (0.3 for 1972-1999 prefecture-year-level), Sasaki (2007) (about 0.2 for 1998-2007 prefecture-quarter-level), and Kambayashi and Ueno (2006) (about 0.3 for 1996-2001 prefecture-month-level). However, the elasticity declines significantly after 2010, reaching close to -0.5 by 2020. The negative elasticity implies that an increase in vacancies results in a decreasing number of matches, which could reflect structural changes in the labor market during COVID-19. The recent sharp decline also highlights a substantial inefficiency in how vacancies are being filled in the labor market.
@@
6.2 Decomposition of part-time and full-time trends in 1972-2023
@@
Given estimation results in each dataset, Figure 4 (e) depicts the trend of implied matching efficiency. The downward trends are common. Full-time matching efficiency fluctuates from 80% to 300%, whereas part-time matching efficiency fluctuates from 60% to 300% compared to January 1972. This concludes that the sharp decline of matching efficiency after COVID-19 shown in the aggregate trend is driven by the decline of matching efficiency for both full-time and part-time workers.
Figure 4 (f) depicts the trends of matching elasticities. The elasticity with respect to unemployment for full-time workers shows moderate fluctuations, particularly between the 1970s and the early 1980s, where it oscillates between 0.4 and 0.8. After the 1980s, this elasticity stabilizes around 0.5. In contrast, the part-time elasticity with respect to unemployment follows a more unstable pattern, fluctuating between 0.1 and 0.9 with significant variations over time. This instability highlights the vulnerability of part-time workers labor markets to economic shocks, particularly during periods of recession, such as the COVID-19 pandemic. The larger fluctuations in elasticity suggest that changes in unemployment among part-time workers have had an uneven and more pronounced impact on the matching process compared to full-time workers.
Figure 4 illustrates some correlation patterns between matching efficiency and market structure variables corresponding with Figures 2 (g)-(h). I find stronger and more consistent relationships in the full-time labor market. This confirms that positive correlations between efficiency and market structure driven mainly by full-time markets.
@@ 6.3 Prefecture-level aggregate results in 2012-2024
Figure 5 displays the month-prefecture level matching efficiency, normalized to 2013 January Tokyo, across different regions of Japan from 2012 to 2024. Each panel shows the efficiency trends within a specific region, highlighting notable regional variations. For instance, the Tohoku region (panel b) and Chubu region (panel d), which are rural areas, exhibit relatively high volatility and generally higher efficiency values compared to other regions, suggesting greater variability in matching efficiency. In contrast, urban regions such as Kanto (panel c) and Kansai (panel e) show more stable and lower efficiency levels. The Tohoku area is about twice as efficient as the Tokyo area. These patterns suggest that regional economic factors and labor market conditions significantly influence matching efficiency across Japan.
@@
6.4 Occupation level results in 2012-2024
Figure 8 illustrates significant disparities in matching efficiency across occupations, normalized to 2013 January for clerical jobs. Primary and secondary industries show relatively higher efficiency than tertiary industries. For example, categories like “Agriculture, Forestry, and Fishing” (panel f) and “Construction and Mining” (panel i) in primary and secondary industries exhibit high volatility and generally higher efficiency levels. In contrast, occupations such as “Managerial” (panel a) and “Service” (panel e) display more stable and lower efficiency values. These variations highlight the diverse nature of labor market dynamics across different sectors, influenced by factors such as skill specialization, economic cycles, and industry-specific demand fluctuations.
Figure 8: Month-occupation level matching efficiency results 2012-2024
Figures 9 and 10 illustrate matching elasticities with respect to unemployed and vacancies in each occupation in 2012-2024. As in the prefecture-level results, there is significant regional heterogeneity within the specific job area. The matching elasticities with respect to unemployed and vacancies fluctuates with a range generally between 0.1 and 6.0, which is larger than prefecture-level results. This indicates that matching heterogeneity depends on the definition of markets and matters in occupation level rather than geographical level.
Figure 9: Month-occupation level matching elasticity with respect to unemployed 2012-2024
Figure 10: Month-occupation level matching elasticity with respect to vacancies 2012-2024